# SPECIAL RELATIVITY

### v9

@@ 6.3 Examination of Stellar aberration in the Context of Special Relativity
It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration discovered by Bradley by claiming it is a consequence of the motion of observers relative to light sources(4).
@@
The lack of symmetry between the cases when either the star or telescope on the earth is moving is shown clearly on the basis of the separation of binary stars. Spectroscopic binaries have velocities exceeding the earth’s velocity round the sun. They revolve around their common center of gravity within days, a period during which the motion of the earth is practically constant. The components of the binary system should be easily separable, when their changing velocities are comparable to the earth’s velocity round the sun. This is, however, not observed (Fig.29-Fig.30) (5).
@@
Figure 29: Aberration shift as inferred from astronomical observations. The plane wave fronts of starlight entering the telescope are imaged by the lens to a diffraction spot which lies in the focal plane. Two cases are selected with different velocities of a earth-based telescope and a star. In the first case (left) the star is at rest with respect to the fixed stars. In the second case (right) the star moves with the same velocity as the earth.
Figure 30: Aberration shift as predicted by the conventional theory of the stellar aberration. Two cases are selected with different velocities of a earth-based telescope and a star. In the first case (left) the star is at rest with respect to the fixed stars. In the second case (right) the star moves with the same velocity as the earth. The conventional theory predicts no aberration of the image in this case.
@@
In Chaper 4, we presented a critical reexamination of the textbook statement that wavefronts and raindrops are to have the same aberration. We used the theory of relativity to show that when one has a transversely moving mirror and a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for light reflected from the mirror. This effect is a consequence of the fact that the Doppler effect is responsible for angular frequency dispersion of light waves reflected from the moving mirror with finite aperture. As a result, the velocity of the energy transport is not equal to the phase velocity. According to the Babinet’s principle, this remarkable prediction of our theory should be correct also for light transmitted through a hole punched in a moving opaque screen or, consequently, through the moving open end of a telescope barrel.
@@
In principle the binary star paradox is resolved by noting that when light passes through the end of a telescope barrel we have a light beam whose fields have been perturbed by diffraction, and, therefore, do not include information about star motion relative to the fixed stars. If the telescope were at rest relative to the fixed stars and the star started to move from rest, then the apparent position of the star seen in the telescope would never jump by any angle.
@@
However, other difficulties arise in the explanation of the earth-based observations, i.e the change in an apparent positions of the fixed stars, which happens when the earth-based telescope changes of its motion relative to the fixed stars. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
@@
An objective of this chapter is to consider non-relativistic interpretations of the stellar aberration, and clearly demonstrate that they are incorrect. Classical kinematics effects leads to serious mistakes if applied to the computation of stellar aberration as seen on the earth rotating around the sun (i.e. in a non-inertial geocentric frame of reference). The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars (6). We derive the aberration for a pulse of light traveling on the surface of the earth using the Langevin metric. Though stellar aberration behaves asymmetrically, it does not contradict special relativity, because the heliocentric (i.e. sun-based) reference system is inertial and the geocentric reference system is non-inertial.
@@ 6.4 Heliocentric Inertial Frame of Reference
Above, we demonstrated that when one has some hole in the opaque screen at rest with respect to the fixed stars and the transversely moving point source, there is no aberration (deviation of the energy transport) for light transmitted through the hole. The absence of the effects of moving (relative to the fixed stars) source in this setup automatically implies the same problem for stellar aberration theory in the heliocentric frame of reference. How shall we solve it? It is like a hole-point source problem with the end of the telescope barrel as a hole.
Suppose that an observer, which is at rest relative to the telescope, performs the direction of the energy transport measurement. At close look at the physics of this subject shows that in the heliocentric frame of reference, where the telescope is at rest, we have actually the problem of steady-state transmission. Then how does the transmitted light beam looks? It looks as though the transmitted beam is going along the telescope axis because it has lost its horizontal group velocity component. That is the transmission appears as shown in Fig 12. It takes the case of a telescope positioned perpendicular to the plane phase front. In other words, the telescope pointed directly at the star. If the motion of the star is parallel to the phase front (i.e. perpendicular to the telescope axis), starlight entering the end of the telescope would be able to pass its full length.
One of the most important conclusions of the discussion presented above is that the aberration of starlight phenomenon is absent in this situation. In particular, the binary components remains unresolved which means that their velocity has no influence on aberration.
@@ 6.5 Earth-Based Non-Inertial Frame of Reference
The analogy between the obliquity of raindrops and the stellar aberration is incorrect. Only on the basis of the theory of relativity and the wave optics, we are able to describe all earth-based experimental observations of stellar aberration. Our theory predicts an effect of stellar aberration in complete agreement to the Bradley’s results, Fig. 31. According to the asymmetry between the inertial and rotating frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the telescope is at rest relative to the earth and the earth rotating relative to the fixed stars, then the direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical sun-based observer. Apparent angle is less than the actual angle. The difference between the actual angle and apparent angle $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=v/c$, where $v$ is the velocity of the earth in its orbit around the sun. It could be said that the crossed term in metric Eq.(15) generates anisotropy in the rotating frame that is responsible for the change of radiation direction (aberration). That is the transmission through the telescope aperture in the rotating frame appears as shown in Fig. 20.
Figure 31: The direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical observer at the sun center. Apparent angle $\phi$ is less than the actual angle $\theta$. The difference between the actual angle and apparent angle is connected with the physical parameters by the relation: $\theta-\phi=v/c$, where $v$ is the velocity of the earth in its orbit around the sun.
The stellar aberration in the geocentric frame of reference is considered independent of the star speed and to have just a local origin exclusively based on the observer speed with respect to the sun.
The main facts which a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems. We presented here a theory which accounts for all these, and in addition gives new results. We demonstrated that the aberration of light is a complex phenomenon which must be branched out into a number of varieties according to their origin. These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (relative to the fixed stars) of the observer or of the light source. Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement. According to our interpretation, there are many kinds of aberration and the stellar aberration in the earth-based frame is only one of these. We demonstrated that all earth-based experiments can be explained on the basis of the effect of the measuring instrument (i.e. physical influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
@@ 6.6 Physical Coordinate System in Space
The aberration of light is the geometric phenomenon. In order to detect the aberration effect inside the earth-based frame, it is obvious that some coordinate system with reference direction is needed. A conventional approach to the aberration of light effect is forcefully based on a definite assumption of reference direction, but this is actually a hidden assumption. Traditionally the physical interpretation of the aberration of light effect in terms of measurements performed with rods that are at rest in the observer’s frame. This (local frame of reference) convention is self-evident and this is the reason why it is never discussed in the aberration of light theory
Fortunately, it is possible to find the space description on a more fundamental level than that of measuring rods. The reference axis in the earth-based frame is formed by gravitation field vector. The plumb-line direction is known as the nadir, leading to the earth’s center. This is the most fundamental local earth-based coordinate system. For example, Bradley used the vertically-mounted telescope. The star he chose was Draconis because it transited almost exactly in zenith. The traditional plumb line provided a sufficiently accurate zenith-point for observations of the stellar aberration.
@@ 6.7 Region of Applicability
The region of applicability of our stellar aberration theory is more wider than one might think. Above we assumed for simplicity that a star image is actually a point spread function in the image plane of a telescope. In other words, it is assumed that the input signal is effectively a plane wave.
@@
Suppose that a telescope is able to distinguish detail in the star image. A complete understanding of the relation between object and image can be obtained if the effect of diffraction are included. The effect of diffraction is to convolve that ideal image with the Fraunhofer diffraction pattern of the telescope pupil.
@@
The star is a spatially completely incoherent source (7). This means that such source is actually a system of elementary (statistically independent) point sources with different offsets. An elementary source with a given offset produces in front of a telescope pupil effectively a plane wave. An elementary source offset tilts the far zone field. Radiation field generated by an completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources. The image of an elementary point source is a point spread function. In other words, there is always the physical influence of the telescope on the measurement of a completely incoherent source. It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like independence on source motion. This argument gives reason why our theory of stellar aberration is correct also for imaging of arbitrary completely incoherent sources.
6.8 Bibliography and Notes
1. The phenomenon of the annual apparent motion of celestial objects about their locations, named stellar aberration, was discovered by Bradley in 1727, who also explained it employing the corpuscular model of light [33].
2. There is a widespread belief that the rays of light coming from the star are falling on the telescope tube and not interact with its sides. To quote French [4]: ”Regarding light as being composed of rain of photons, we can easily calculate the change in apparent direction of a distant object such a star.” Similar statements can also be found in recently published textbooks. To quote Rafelski [5]: ”We consider a light ray originating in a distant star. … The following discussion is addressing the observation of well focusing light-rays and not (spherical) plane wave light. … This picture of the experimental situation is accurate since the light emitted by a star consists of an incoherent flux of photons produced in independent atomic processes.” This conclusion is incorrect. We emphasize that the design of a telescope optical system is based on the classical theory of diffraction (originated directly from the classical electromagnetic theory). The resolution of a telescope is limited by the diffraction of light waves and the telescope always influences the measured radiation (due to unavoidable diffraction of a star light by the telescope aperture). In a well corrected optical system and in a circular pupil, the size Airy disc of diffraction pattern (i.e. size of the image of a point source) is inversely proportional to the diameter of the pupil.
3. A star is a completely incoherent source. The character of the mutual intensity function produced by an incoherent source is fully described by the Van Cittert-Zernike theorem [34]. Any star can be considered as very far away from the sun. In all cases of practical interest, telescopes are situated in the far-zone of the (distant star) source. Consider a star like Sirius, which is one of the the nearest stars. The coherent area of light observed from Sirius has a diameter of about 6 m. This correlation was observed by Brown and Twiss in 1956 [35].
4. It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer. In the paper on the theory of relativity Einstein deduced the aberration formula from the idea that the velocity of $v$ is the relative velocity of the star-earth system. The idea was represented by many authors of textbooks. To quote Moeller [13]: ”This phenomenon, which is called aberration, was observed … by Bradley who noticed that the stars seem to perform a collective annual motion in the sky. This apparent motion is simply due to the fact that the observed direction of a light ray coming from a star depends on the velocity of the earth relative to the star.”
5. In 1950 Ives [36] stressed for the first time that the presence of binaries in the sky gave rise to an important difficulty for the theory of relativity. It is stated that the idea that aberration may be described in terms of relative motions of the bodies concerned is immediately refuted by the existence of spectroscopic binaries with velocities comparable with that of the Earth in its orbit. Still this exhibit aberrations not different from other stars. For example, a spectroscopic binary, Mizar A, has well-known orbital parameters, from which can be calculated an observable angular separation of 1’10” if aberration were due to relative velocity. The empirical value is less than 0.01”, clearly incompatible with authors of textbooks point of view [37, 38]. There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions.
6. Stellar aberration exist as observable phenomenon only in the presence of changing states of motion (i.e. acceleration). The problem of the earth-based measurements is related with the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars. This has been recognized by some expects, perhaps most explicitly by Selleri who states [38] that ”Thus a complete explanation of the aberration effect is given in terms of variations of the earth absolute velocity due to orbital motion, while the star/earth relative velocity is irrelevant. Thus acceleration (of planet, this time) enters once more into a game.”
7. From everything that authors of textbooks have written about the star light (”rain of photons”) it clearly follows that they did not understand the several aspects of the statistical optics. In the statistical optics for star emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as a ensemble of atoms) by quantum mechanics. The atomic polarization induced by the radiation field appears as a driving term in Maxwell’s equations and sustains the oscillations. The star is actually a system of elementary (statistically independent) point sources with different offsets. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. Semi-classical theory is treating these atoms as coherent, interacting radiating dipoles. The induced macroscopic dipole moment in an elementary source leads to the classical electromagnetic radiation. The distinguishing characteristic of the statistical optics is the fact that electromagnetic fields are treated in a completely classical manner until they interact with the atoms of the photosensitive material on which they are incident. Thus there is no necessity to deal with quantization of the electromagnetic field; only the interaction of the classical field and matter is quantized. Thus in the semiclassical theory, where the atom is treated quantum mechanically but the electromagnetic field is treated classically, the photodetector has the effect of converting the continuous cycle-averaged classical intensity of the light beam into a succession of discrete photocounts. Light interact with matter in a fundamentally random way. As a consequence, any measurement of (star) light will be accompanied by certain unavoidable fluctuations in the detector. Stochastic fluctuations of the classical intensity can influence the statistical properties of the photoevents that are observed. Note in particular that the variance $\sigma^{2}_{K}$ of photoevents $K$ consists of two distinct terms, each of which has a physical interpretation: $\sigma^{2}_{K}=[K^{2}-<K>^{2}]/<K>^{2}=1/<K>+\sigma^{2}_{W}$. The first term is simply the variation of the counts that would be observed if classical intensities were constant and the photocounts were purely Poisson. We refer to this contribution to the count fluctuations as ”quantum noise”. The second therm, $\sigma^{2}_{W}$, is clearly zero if there are no fluctuations of the classical intensity. For instance, in the case of laser light, this component would be identically zero, and the count variance would be simply that arising from the photon ”shot noise”. This is intrinsic fluctuations associated with quantum fluctuations of vacuum and can give information only on the mean rate ot photons production. When star (thermal) light is incident on the photodetector, the classical fluctuations are nonzero, and the variance of the photocounts is larger than that expected for a Poisson distribution. The classical thermal fluctuations contains rich information about radiation source. In particular, thermal noise provide us with information for analyzing (using Van Cittert-Zernike theorem) the spatial source profile. The ratio of the classical variance to the ”photon shot noise” variance is equal to $\delta_{c}=<K>\sigma^{2}_{W}$. Parameter $\delta_{c}$ named as the photocount degeneracy parameter and can be interpreted as the average number of counts that occur in a single coherence interval of the incident radiation. When $\delta_{c}\ll 1$, it is highly probable that will be no more than one count per coherence interval of the wave, with the result that shot noise predominates over classically induced noise. On the other hand, when $\delta_{c}\gg 1$, there are many photoevents present in each coherence interval of the wave. The result is ”bunching” of the photoevents by the classical intensity fluctuations and an increase of the variance of the counts to the point where the classically induced fluctuations are far stronger than the shot noise variations.

### v10

@@ 6.3 Examination of Stellar aberration in the Context of Special Relativity
It is generally believed that the theory of relativity appears to conform to the phenomenon of stellar aberration discovered by Bradley by claiming it is a consequence of the motion of observers relative to light sources 313131 In the paper on the theory of relativity, Einstein deduced the aberration formula from the idea that the velocity of $v$ is the relative velocity of the star-earth system. The idea was represented by many authors of textbooks. To quote Moeller [13]: ”This phenomenon, which is called aberration, was observed … by Bradley who noticed that the stars seem to perform a collective annual motion in the sky. This apparent motion is simply because the observed direction of a light ray coming from a star depends on the velocity of the earth relative to the star.”.
@@
The lack of symmetry between the cases when either the star or telescope on the earth is moving is shown clearly based on the separation of binary stars. Spectroscopic binaries have velocities exceeding the earth’s velocity around the sun. They revolve around their common center of gravity within days, a period during which the motion of the earth is practically constant. The components of the binary system should be easily separable when their changing velocities are comparable to the earth’s velocity around the sun. This is, however, not observed (Fig.27-Fig.28)  [^32] In 1950 Ives [36] stressed for the first time that the presence of binaries in the sky gave rise to an important difficulty for the theory of relativity. It is stated that the idea that aberration may be described in terms of relative motions of the bodies concerned is immediately refuted by the existence of spectroscopic binaries with velocities comparable with that of the Earth in its orbit. Still, this exhibit’s aberrations are not different from other stars. For example, a spectroscopic binary, Mizar A, has well-known orbital parameters, from which can be calculated an observable angular separation of 1’10” if aberration were due to relative velocity. The empirical value is less than 0.01”, clearly incompatible with authors of textbooks point of view [37, 38]..
@@
Figure 27: Aberration shift as inferred from astronomical observations. The plane wavefronts of starlight entering the telescope are imaged by the lens to a diffraction spot that lies in the focal plane. Two cases are selected with different velocities of an earth-based telescope and a star. In the first case (left) the star is at rest with respect to the fixed stars. In the second case (right) the star moves with the same velocity as the earth.
Figure 28: Aberration shift as predicted by the conventional theory of the stellar aberration.
@@
In Chapter 4, we presented a critical reexamination of the textbook statement that wavefronts and raindrops are to have the same aberration. We used the theory of relativity to show that when one has a transversely moving mirror and a plane wave of light is falling normally on the mirror, there is a deviation of the energy transport for light reflected from the mirror. This effect is a consequence of the fact that in the absolute time coordinatization, the wave equation will depend on the velocity vector. As a result, the velocity of the energy transport is not equal to the phase velocity. According to Babinet’s principle, this remarkable prediction of our theory should be correct also for light transmitted through a hole punched in a moving opaque screen or, consequently, through the moving open end of a telescope barrel.
@@
In principle, the binary star paradox is resolved by noting that when light passes through the end of a telescope barrel we have a light beam whose fields have been perturbed by diffraction, and, therefore, do not include information about star motion relative to the fixed stars. If the telescope were at rest relative to the fixed stars and the star started to move from rest, then the apparent position of the star seen in the telescope would never jump by any angle.
@@
However, other difficulties arise in the explanation of the earth-based observations, i.e. the change in apparent positions of the fixed stars, which happens when the earth-based telescope changes its motion relative to the fixed stars. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration.
@@
The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars  [^33] This has been recognized by some experts, perhaps most explicitly by Selleri who states [38] that ”Thus a complete explanation of the aberration effect is given in terms of variations of the earth absolute velocity due to orbital motion, while the star/earth relative velocity is irrelevant. Thus acceleration (of the planet, this time) enters once more into a game.”. We derive the aberration for a pulse of light traveling on the surface of the earth using the Langevin metric. Even though stellar aberration behaves asymmetrically, it does not contradict special relativity, because the heliocentric (i.e. sun-based) reference system is inertial and the geocentric reference system is non-inertial.
@@ 6.4 Heliocentric Inertial Frame of Reference
Above, we demonstrated that when one has some hole in the opaque screen at rest with respect to the fixed stars and the (tangentially to the screen) moving point source, there is no aberration (deviation of the energy transport) for light transmitted through the hole. The transmission appears as shown in Fig 12. The absence of the effects of moving (relative to the fixed stars) source in this setup automatically implies the same problem for stellar aberration theory in the heliocentric frame of reference. How shall we solve it? It is like a hole-point source problem with the end of the telescope barrel as a hole.
Suppose that an observer, which is at rest relative to the telescope, performs the direction of the energy transport measurement. A close look at the physics of this subject shows that in the heliocentric frame of reference, where the telescope is at rest, we have the problem of steady-state transmission. Then how does the transmitted light beam look? It looks as though the transmitted beam is going along the telescope axis because it has lost its horizontal group velocity component. It takes the case of a telescope positioned perpendicular to the plane phase front. In other words, the telescope pointed directly at the star. If star’s motion is parallel to the phase front (i.e. perpendicular to the telescope axis), starlight entering the end of the telescope would be able to pass its full length.
One of the most important conclusions of the discussion presented above is that the aberration of the starlight phenomenon is absent in this situation. In particular, the binary components remain unresolved which means that their velocity does not influence aberration.
@@ 6.5 Earth-Based Non-Inertial Frame of Reference
The analogy between the obliquity of raindrops and the stellar aberration is incorrect. Only based on the theory of relativity and wave optics, we are able to describe all earth-based experimental observations of stellar aberration. Our theory predicts an effect of stellar aberration in complete agreement with Bradley’s results, Fig. 29. According to the asymmetry between the inertial and rotating frames, there is a remarkable prediction on the theory of the aberration of light. Namely, if the telescope is at rest relative to the earth and the earth is rotating relative to the fixed stars, then the direction of a star as seen from the earth is not the same as the direction when viewed by a hypothetical sun-based observer. The apparent angle is less than the actual angle. The difference between the actual angle and apparent angle $\theta_{a}$ is connected with the physical parameters by the relation: $\theta_{a}=v/c$, where $v$ is the velocity of the earth in its orbit around the sun. It could be said that the crossed term in metric Eq.(15) generates anisotropy in the rotating frame that is responsible for the change of radiation direction (aberration). The transmission through the telescope aperture in the rotating frame appears as shown in Fig.18 - Fig. 19.
Figure 29: The direction of a star as seen from the earth.
The stellar aberration in the geocentric frame of reference is considered independent of the star’s speed and to have just a local origin exclusively based on the observer’s speed with respect to the sun. According to our interpretation, there are many kinds of aberration and the stellar aberration in the earth-based frame is only one of these. We demonstrated that all earth-based experiments can be explained based on the effect of the measuring instrument (i.e. physical influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
@@ 6.6 Physical Coordinate System in Space
The aberration of light is a geometric phenomenon. To detect the aberration effect inside the earth-based frame, it is obvious that some coordinate system with reference direction is needed. A conventional approach to the aberration of light effect is forcefully based on a definite assumption of reference direction, but this is a hidden assumption. Traditionally the physical interpretation of the aberration of light effect in terms of measurements performed with rods that are at rest in the observer’s frame. This (local frame of reference) convention is self-evident and this is the reason why it is never discussed in the aberration of light theory
Fortunately, it is possible to find the space description on a more fundamental level than that of measuring rods. The reference axis in the earth-based frame is formed by a gravitation field vector. The plumb-line direction is known as the nadir, leading to the earth’s center. This is the most fundamental local earth-based coordinate system. For example, Bradley used the vertically-mounted telescope. The star he chose was Draconis because it transited almost exactly in zenith. The traditional plumb line provided a sufficiently accurate zenith-point for observations of the stellar aberration.
@@ 6.7 Region of Applicability
The region of applicability of our stellar aberration theory is more wider than one might think. Above we assumed for simplicity that a star image is a point spread function in the image plane of a telescope. In other words, it is assumed that the input signal is effectively a plane wave.
@@
Suppose that a telescope is able to distinguish detail in the star image. A complete understanding of the relation between object and image can be obtained if the effect of diffraction is included. The effect of diffraction is to convolve that ideal image with the Fraunhofer diffraction pattern of the telescope pupil.
@@
The star is a spatially completely incoherent source  [^34] In the statistical optics for star emission, the fields are described classically at the level of Maxwell’s equations and the emitting medium is treated (as an ensemble of atoms) by quantum mechanics. The star is a system of elementary point sources. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength. Within an elementary source volume ($\lambda^{3}$) there is an enormous number of atoms. The semi-classical theory is treating these atoms as coherent, interacting radiating dipoles. The induced macroscopic dipole moment in an elementary source leads to classical electromagnetic radiation. The distinguishing characteristic of statistical optics is the fact that electromagnetic fields are treated in a completely classical manner until they interact with the atoms of the photosensitive material on which they are incident. Thus there is no necessity to deal with quantization of the electromagnetic field; only the interaction of the classical field and matter is quantized. The photodetector has the effect of converting the continuous cycle-averaged classical intensity of the light beam into a succession of discrete photocounts.. This means that such a source is a system of elementary (statistically independent) point sources with different offsets. An elementary source with a given offset produces in front of a telescope pupil effectively a plane wave. An elementary source offset tilts the far zone field. A radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individual elementary point sources. The image of an elementary point source is a point spread function. In other words, there is always the physical influence of the telescope on the measurement of a completely incoherent source. It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like independence on source motion. This argument gives the reason why our theory of stellar aberration is correct also for imaging of arbitrary completely incoherent sources.